# Mixed Hodge Modules and Canonical Perverse Extensions for Multi-Node Conifold Degenerations

### 07 Apr 2026 v1

@@ 6.3. The global gluing datum
The assembled datum $(M^{\prime}_{U},M^{\prime\prime}_{\Sigma},u_{\Sigma},v_{\Sigma})$ satisfies the precise compatibility required for Saito’s principal-divisor gluing theorem.
@@
By Lemma 6.1, the singular quotient object is $\mathcal{M}^{\prime\prime}_{\Sigma}$. By Proposition 6.2, the local gluing maps assemble to global morphisms $u_{\Sigma}$ and $v_{\Sigma}$. By Proposition 6.3, these morphisms satisfy the required relation
$v_{\Sigma}u_{\Sigma}=N$
on the finite node-supported ODP sector. Therefore the data satisfy the hypotheses of Saito’s principal-divisor gluing theorem. ∎
@@ 6.4. Construction of the global corrected mixed Hodge module
whose restriction to the smooth locus agrees with $IC^{H}_{X_{0}}|_{U}=\mathbb{Q}^{H}_{U}[3]$ and whose singular quotient is
@@
Apply Saito’s principal-divisor gluing theorem to the global gluing datum of Theorem 6.4. This produces an object
@@
with open-part restriction $\mathcal{M}^{\prime}_{U}=\mathbb{Q}^{H}_{U}[3]$ and singular quotient
@@
The canonical middle extension of $\mathbb{Q}^{H}_{U}[3]$ across $X_{0}$ is $IC^{H}_{X_{0}}$, so $\mathcal{P}^{H}$ is the desired global corrected mixed Hodge module. ∎
@@ 6.8. Global uniqueness and rigidity
    $\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1);$
  3. (3)
    for each node $p_{k}\in\Sigma$, its restriction to a sufficiently small analytic neighborhood of $p_{k}$ is isomorphic to the local mixed-Hodge-module ODP extension of Theorem 4.11 (equivalently, the unique local object of Proposition 4.14).
@@
Both $\mathcal{E}^{H}$ and $\mathcal{P}^{H}$ determine extension classes in
@@
By Lemma 6.10, each such global class is completely determined by its nodewise components in the direct-sum decomposition
@@
Fix a node $p_{k}\in\Sigma$. By assumption, the restriction of $\mathcal{E}^{H}$ to a sufficiently small analytic neighborhood of $p_{k}$ is isomorphic to the local mixed-Hodge-module ODP extension of Theorem 4.11. On the other hand, the global object $\mathcal{P}^{H}$ was constructed by assembling exactly these local ODP blocks via the global gluing procedure of Section 6, so its restriction near $p_{k}$ is the same local object. By the local rigidity statement of Proposition 4.14, the corresponding local extension class is uniquely determined. Therefore the $k$-th nodewise component of the global extension class of $\mathcal{E}^{H}$ agrees with the $k$-th nodewise component of the global extension class of $\mathcal{P}^{H}$.
Since this holds for every node $p_{k}\in\Sigma$, all nodewise components of the two global classes agree. Lemma 6.10 now implies that the two global extension classes in
$\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{M}^{\prime\prime}_{\Sigma},IC^{H}_{X_{0}})$
are equal. Hence the nodewise components of the extension classes of $\mathcal{E}^{H}$ and $\mathcal{P}^{H}$ agree at every node, so by Lemma 6.10 the global extension classes agree, and therefore $\mathcal{E}^{H}\cong\mathcal{P}^{H}$. ∎
@@ 7.3. Hypercohomology / LMHS comparison theorem
The mixed-Hodge-module extension in Eqn. (7.1) first produces an extension on hypercohomology in the category of mixed Hodge structures, and the claim is that this induced extension is exactly the finite vanishing-sector extension coming from the LMHS nearby-cycle formalism. By Proposition 6.9,
@@
Thus $\mathcal{P}^{H}$ is the global object produced from the same nearby-cycle data that carry the LMHS.
@@
The top row is obtained from Eqn. (7.1). The bottom row is obtained from the nearby-cycle formalism after restricting to the finite point-supported vanishing contribution $\mathcal{V}^{H}$. Since $\mathcal{P}^{H}$ is the global corrected object produced from the same nearby-cycle and variation data, and
$\mathcal{P}^{H}\simeq\operatorname{Cone}(v_{\Sigma})[-1]$
by Proposition 6.9, the two constructions come from the same distinguished triangle in the derived category of mixed Hodge modules. Therefore the associated long exact sequences agree, yielding the commutative diagram. ∎
@@
By Lemma 7.5, the connecting morphisms in the long exact sequence (7.2) coincide with the connecting morphisms in the nearby-cycle long exact sequence on the finite vanishing sector. By Definition 7.4, the latter determine the LMHS extension class. Therefore the extension induced by (7.1) on hypercohomology is canonically the LMHS extension class. ∎
@@ 7.4. Extension classes and comparison questions
This is the standard compatibility of a cohomological functor with Yoneda extensions in the heart of a $t$-structure: applying hypercohomology to Eqn. (7.1) yields the long exact sequence Eqn. (7.2), and the connecting morphisms encode the induced extension class in mixed Hodge structures. ∎
@@
The first map is induced by exactness of $\mathrm{rat}$. The second is induced by hypercohomology together with Lemma 7.7. By Theorem 7.6, the induced hypercohomological extension is canonically the LMHS extension class. ∎
@@
The relation between the first two classes is given by exactness of $\mathrm{rat}$. The relation between the first and third is given by Lemma 7.7 and Theorem 7.6. The final assertion follows from the construction of $\mathcal{P}^{H}$ from the nearby-cycle and variation data in Sections 4–6. ∎
@@ 8.4. Toward the domain-wall interpretation
The significance of the present paper is that this interpretation no longer rests only on the perverse side. Theorems 6.8 and 7.6 show that the same localized quotient also contributes the vanishing part of the limiting mixed Hodge structure on hypercohomology. Thus the bulk/localized-sector interpretation is simultaneously visible in the mixed-Hodge-module extension, in its perverse realization, and in the Hodge-theoretic degeneration data. This is the precise mathematical content that later physical applications may safely use.
@@ 9.1. What has been proved
The main result of the present paper is the construction of a fully internal mixed-Hodge-module refinement of the canonical corrected perverse object for a finite multi-node conifold degeneration. More precisely, for a one-parameter degeneration whose central fiber $X_{0}$ has ordinary double points
$\Sigma=\{p_{1},\dots,p_{r}\},$
we constructed an object
@@
whose realization is the corrected perverse object
$\mathcal{P}.$
This gives a theorem-level Hodge-theoretic refinement of the canonical corrected perverse extension in the finite multi-node ordinary double point setting.
The construction proceeds by combining four essential ingredients. First, the local ordinary double point model yields a rank-one point-supported mixed-Hodge-module block at each singular point. Second, the finite support decomposition identifies the global singular contribution as the direct sum of these local rank-one blocks. Third, Saito’s divisor-case gluing formalism provides the mechanism for assembling the smooth-locus Hodge-module data and the finite node contributions into a single global object on $X_{0}$. Fourth, realization carries the resulting mixed-Hodge-module extension to the canonical corrected perverse object, while hypercohomology identifies the same point-supported quotient with the vanishing sector of the limiting mixed Hodge structure.
The resulting theorem package may be summarized as follows:
  • there exists an object
    $\mathcal{P}^{H}\in MHM(X_{0});$
  • its realization satisfies
    $\mathrm{rat}(\mathcal{P}^{H})\cong\mathcal{P};$
  • it fits into an exact sequence
    in $MHM(X_{0})$;
  • the point-supported quotient realizes the rank-$r$ local vanishing contribution of the degeneration;
  • the hypercohomology of $\mathcal{P}^{H}$ is functorially related to the limiting mixed Hodge structure through the same nearby-cycle mixed-Hodge-module formalism;
  • and the corrected extension is organized simultaneously as a class in mixed Hodge modules, as its realized class in perverse sheaves, and as an induced extension class on hypercohomology in mixed Hodge structures.
In particular, the present paper upgrades the corrected perverse extension from a constructible or perverse object to a genuine mixed-Hodge-module object carrying internal Hodge-theoretic structure. This is precisely the refinement that was identified as missing in the earlier nearby-cycle bridge picture and that was not available in the original perverse-sheaf construction.
@@ 9.2. Consequences for subsequent work
The significance of the present paper is not only that it proves a new mixed-Hodge-module theorem, but that it changes the status of the larger program built around the corrected perverse object. Before the present construction, one could compare the corrected perverse extension and the Hodge-theoretic degeneration data only at the level of a common nearby-cycle origin. After the present paper, one has an actual object
@@
whose realization is the corrected perverse extension and whose point-supported quotient is the Hodge-theoretic refinement of the local vanishing sector.
This has several immediate consequences.
First, it supplies a theorem-level foundation for the study of global extension data in finite multi-node degenerations. The corrected extension is no longer merely an object of $\mathrm{Perv}(X_{0};\mathbb{Q})$; it is now an extension in the category of mixed Hodge modules, with an internal Hodge-theoretic quotient and an induced extension on hypercohomology. Thus later work may safely treat the global extension data as genuinely Hodge-theoretic rather than merely perverse.
Second, the extension space itself is now substantially better understood. The singular quotient is the finite direct sum
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),$
and the corresponding global extension class decomposes nodewise. On the perverse side, the finite ordinary double point setting yields distinguished nodewise extension classes
$e_{1},\dots,e_{r},$
so that the corrected perverse extension admits a concrete nodewise organization. This is the first precise mathematical form of the statement that the degeneration contains finitely many localized sectors coupled to a bulk geometric sector.
Third, it gives a rigorous mathematical substrate for the quiver-theoretic, schober-theoretic, and wall-crossing directions anticipated earlier. The finite collection of rank-one point-supported objects
$\{i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\}_{k=1}^{r}$
together with the global extension class of $\mathcal{P}^{H}$ already has the formal shape of a finite system of localized sectors coupled to a bulk geometric sector. This is exactly the kind of structure from which quiver, schober, and wall-crossing formalisms are expected to emerge.
Fourth, it provides theorem-level mathematical support for subsequent physical interpretations of multi-node conifold degenerations. In particular, the exact sequence
gives a precise mixed-Hodge-module version of the bulk/localized-sector decomposition that later work may interpret in physical language. The important point is that this interpretation no longer rests on heuristic analogy alone: the localized quotient is now an actual theorem-level object in $MHM(X_{0})$, and its contribution to the limiting mixed Hodge structure is part of the proved hypercohomological picture.
In this sense, the present paper should be viewed as the technical keystone between the earlier perverse-sheaf construction and the later categorical and physical developments. It does not yet solve the full wall-crossing or schober problem, but it supplies the foundational Hodge-theoretic object that those later theories require.

### 09 Apr 2026 v2

@@ 6.3. The global gluing datum
We now package the results of Sections 6.1 and 6.2 into the form required by Saito’s principal-divisor gluing theorem.
@@
Lemma 6.1 identifies the singular term as
$\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),$
and Proposition 6.2 provides the assembled global attaching maps $u_{\Sigma}$ and $v_{\Sigma}$. By Proposition 6.4, these data satisfy the Hodge, weight, and Kashiwara–Malgrange $V$-filtration compatibilities required by Saito’s principal-divisor gluing theorem, together with the compatibility relation
$v_{\Sigma}u_{\Sigma}=N.$
Therefore
is a valid divisor-gluing datum in the sense of [12, Prop. 0.3]. ∎
@@ 6.4. Construction of the global corrected mixed Hodge module
Applying Saito’s gluing theorem to the datum of Theorem 6.5 produces the global mixed Hodge module refining the corrected perverse object.
@@
whose restriction to the smooth locus agrees with
$IC^{H}_{X_{0}}\big|_{U}\cong\mathbb{Q}^{H}_{U}[3]$
and whose singular quotient is
@@
By Theorem 6.5 and Proposition 6.4, the global ordinary-double-point package defines an admissible filtered gluing datum in the sense of Saito’s divisor-case gluing theorem [12, Prop. 0.3]. Applying that theorem yields an object
@@
whose restriction to the smooth locus is $\mathbb{Q}^{H}_{U}[3]$ and whose singular quotient is
@@
The canonical middle extension of $\mathbb{Q}^{H}_{U}[3]$ from $U$ to $X_{0}$ is $IC^{H}_{X_{0}}$. Hence the glued object fits into a short exact sequence
$0\to IC^{H}_{X_{0}}\to\mathcal{P}^{H}\to\mathcal{M}^{\prime\prime}_{\Sigma}\to 0.$
This is exactly the required global corrected mixed Hodge module. ∎
@@ 6.8. Global uniqueness and rigidity
Proposition 6.12 (Recovery of the local ODP model from the global construction).
Let $\mathcal{P}^{H}$ be the global corrected mixed Hodge module of Theorem $6.$ For each node $p_{k}\in\Sigma$, the restriction of $\mathcal{P}^{H}$ to a sufficiently small analytic neighborhood of $p_{k}$ is isomorphic to the local mixed-Hodge-module ODP extension of Theorem 4.13.
Proof.
Fix a node $p_{k}\in\Sigma$, and choose a sufficiently small analytic neighborhood $B_{k}\subset X_{0}$ containing no singular point other than $p_{k}$. By construction, the global singular quotient
$\mathcal{M}^{\prime\prime}_{\Sigma}=\bigoplus_{j=1}^{r}i_{j*}\mathbb{Q}^{H}_{\{p_{j}\}}(-1)$
restricts on $B_{k}$ to the single local summand
$i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Likewise, the global attaching maps $u_{\Sigma}$ and $v_{\Sigma}$ restrict on $B_{k}$ to the local attaching maps $u_{k}$ and $v_{k}$, because the global datum was assembled nodewise from the local ODP blocks and all other summands vanish on $B_{k}$.
Therefore the restriction of the global gluing datum to $B_{k}$ is exactly the local gluing datum of Theorem 4.10. Applying Saito’s gluing theorem locally on $B_{k}$, the restricted object $\mathcal{P}^{H}|_{B_{k}}$ is the mixed Hodge module produced from that local datum. By Proposition 4.16, that local object is unique up to unique isomorphism, hence
$\mathcal{P}^{H}|_{B_{k}}\cong\mathcal{P}^{H}_{\mathrm{loc},k}.$
This is exactly the asserted local recovery statement. ∎
@@
Both $\mathcal{E}^{H}$ and $\mathcal{P}^{H}$ define extension classes in
@@
By Lemma 6.11, each such global class is determined by its nodewise components under the decomposition
@@
Fix a node $p_{k}\in\Sigma$, and choose a sufficiently small analytic neighborhood $B_{k}\subset X_{0}$ containing no other node. Restriction is exact, so $\mathcal{E}^{H}|_{B_{k}}$ fits into a short exact sequence
$0\to IC^{H}_{X_{0}}|_{B_{k}}\to\mathcal{E}^{H}|_{B_{k}}\to i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)\to 0.$
Moreover,
$\mathrm{rat}(\mathcal{E}^{H}|_{B_{k}})\cong\mathrm{rat}(\mathcal{E}^{H})|_{B_{k}}\cong\mathcal{P}|_{B_{k}},$
and $\mathcal{P}|_{B_{k}}$ is the corrected local perverse ODP extension. Thus $\mathcal{E}^{H}|_{B_{k}}$ has the same realized local corrected perverse block and the same point-supported singular quotient as the local object $\mathcal{P}^{H}_{\mathrm{loc},k}$.
By Proposition 4.16, the local mixed-Hodge-module extension with these properties is unique up to isomorphism. Therefore
$\mathcal{E}^{H}|_{B_{k}}\cong\mathcal{P}^{H}_{\mathrm{loc},k}.$
By Proposition 6.12, the same holds for $\mathcal{P}^{H}|_{B_{k}}$. Hence the $k$-th nodewise extension component of $\mathcal{E}^{H}$ agrees with that of $\mathcal{P}^{H}$.
Since this holds for every node $p_{k}\in\Sigma$, all nodewise components agree. Lemma 6.11 then implies that the global extension classes of $\mathcal{E}^{H}$ and $\mathcal{P}^{H}$ coincide in
$\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{M}^{\prime\prime}_{\Sigma},IC^{H}_{X_{0}}),$
and therefore
$\mathcal{E}^{H}\cong\mathcal{P}^{H}.$
@@ 7.3. Hypercohomology / LMHS comparison theorem
The point of this subsection is to identify the extension carried by Eqn. (7.1) after passing to hypercohomology with the finite vanishing-sector extension arising from the nearby-cycle description of the limiting mixed Hodge structure.
By Proposition 6.10,
@@
Thus $\mathcal{P}^{H}$ is not an auxiliary object introduced after the fact: it is the global corrected object canonically constructed from the same nearby-cycle and variation data that govern the limiting mixed Hodge structure.
@@
Proposition 7.5.
There is a distinguished triangle in $D^{b}MHM(X_{0})$
$IC^{H}_{X_{0}}\longrightarrow\mathcal{P}^{H}\longrightarrow\mathcal{V}^{H}\overset{+1}{\longrightarrow}$
whose hypercohomology long exact sequence is Eqn. (7.2). Moreover, this triangle is the finite vanishing-sector truncation of the nearby-cycle triangle determined by the same global variation morphism $v_{\Sigma}$.
Proof.
The exact sequence (7.1) is an exact sequence in the abelian category $MHM(X_{0})$, hence determines a distinguished triangle
$IC^{H}_{X_{0}}\longrightarrow\mathcal{P}^{H}\longrightarrow\mathcal{V}^{H}\overset{+1}{\longrightarrow}$
in $D^{b}MHM(X_{0})$. Applying hypercohomology yields the long exact sequence (7.2).
For the second assertion, Proposition 6.10 identifies $\mathcal{P}^{H}$ with $\operatorname{Cone}(v_{\Sigma})[-1]$, where $v_{\Sigma}$ is the global morphism built from the nearby-cycle and variation data along the finite set of nodes $\Sigma$. By Theorem 6.9, the quotient of $\mathcal{P}^{H}$ by $IC^{H}_{X_{0}}$ is precisely
$\mathcal{V}^{H}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),$
so the above distinguished triangle is exactly the triangle obtained by isolating the finite point-supported vanishing contribution inside the nearby-cycle construction. In this sense it is the finite vanishing-sector truncation of the nearby-cycle triangle. ∎
@@
The top row is the long exact sequence attached to the distinguished triangle of Proposition 7.5. For the bottom row, start with the nearby-cycle triangle for $\psi_{\pi}^{H}(F)$ and then pass to the finite point-supported vanishing summand identified by
$\mathcal{V}^{H}=\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1).$
Because Proposition 6.10 identifies $\mathcal{P}^{H}$ with the cone of the same global variation morphism $v_{\Sigma}$, both rows are obtained functorially from the same nearby-cycle/variation data after isolating the finite vanishing contribution. The vertical maps are the evident identifications on the intersection-complex term and on the point-supported vanishing term, together with the induced comparison morphism on the middle term. Naturality of the long exact sequence associated to a distinguished triangle then gives the claimed commutative diagram. ∎
@@
By Proposition 7.5, the exact sequence (7.1) determines a distinguished triangle whose hypercohomology long exact sequence is precisely (7.2). By Lemma 7.6, this long exact sequence agrees with the nearby-cycle long exact sequence after restriction to the finite point-supported vanishing sector. In particular, the connecting morphisms
$\partial_{m}:\mathbb{H}^{m}(X_{0},\mathcal{V}^{H})\longrightarrow\mathbb{H}^{m+1}(X_{0},IC^{H}_{X_{0}})$
are exactly the connecting morphisms that define the finite vanishing-sector extension on the limiting mixed Hodge structure.
By Definition 7.4, the LMHS extension class is, by definition, the extension class in $MHS$ encoded by these connecting morphisms on the finite vanishing sector. Hence the extension induced by (7.1) on hypercohomology is canonically the LMHS extension class. The equivalent formulation follows immediately, since $\mathcal{V}^{H}$ is precisely the point-supported quotient term appearing in (7.1). ∎
@@ 7.4. Extension classes and comparison questions
Let
$0\longrightarrow IC^{H}_{X_{0}}\longrightarrow\mathcal{P}^{H}\longrightarrow\mathcal{V}^{H}\longrightarrow 0$
be the exact sequence (7.1) in the abelian category $MHM(X_{0})$. Any short exact sequence in the heart of a $t$-structure determines both a Yoneda class in $\mathrm{Ext}^{1}$ and a distinguished triangle in the derived category. Applying the cohomological functor $\mathbb{H}^{\bullet}(X_{0},-)$ to this distinguished triangle yields the long exact sequence (7.2).
The connecting morphisms in that long exact sequence are functorially determined by the original short exact sequence, hence by its Yoneda class $\mathcal{P}^{H}]$. Equivalently, the image of $\mathcal{P}^{H}]$ under the extension map induced by hypercohomology is the class represented by the connecting morphisms of (7.2). Therefore the hypercohomology functor sends the Yoneda class $\mathcal{P}^{H}]$ to the extension class in $MHS$ determined by those connecting morphisms. ∎
@@
Because $\mathrm{rat}:MHM(X_{0})\to\mathrm{Perv}(X_{0};\mathbb{Q})$ is exact, it sends short exact sequences in $MHM(X_{0})$ to short exact sequences in $\mathrm{Perv}(X_{0};\mathbb{Q})$, and therefore induces the first map on $\mathrm{Ext}^{1}$. Applying $\mathrm{rat}$ to Eqn. (7.1) gives the perverse exact sequence (7.3), so the class $\mathcal{P}^{H}]$ maps to the realized perverse extension class $\mathcal{P}]$.
For the second map, the cohomological functor $\mathbb{H}^{\bullet}(X_{0},-)$ sends the short exact sequence (7.1) to the long exact sequence (7.2); by Lemma 7.8, the image of $\mathcal{P}^{H}]$ is the extension class in $MHS$ encoded by the connecting morphisms of that long exact sequence. By Theorem 7.7, this class is canonically the LMHS extension class. This proves both assertions. ∎
@@
The mixed-Hodge-module extension class in (1) is, by definition, the Yoneda class
$\mathcal{P}^{H}]\in\mathrm{Ext}^{1}_{MHM(X_{0})}(\mathcal{V}^{H},IC^{H}_{X_{0}})$
of the exact sequence (7.1). Since the realization functor is exact, applying $\mathrm{rat}$ to (7.1) yields the perverse exact sequence (7.3); this identifies the class in (2) as the functorial image of the class in (1).
Next, applying hypercohomology to (7.1) yields the long exact sequence (7.2). By Lemma 7.8, the Yoneda class $\mathcal{P}^{H}]$ maps to the extension class in $MHS$ determined by the connecting morphisms of that long exact sequence. By Theorem 7.7, this induced class is exactly the LMHS extension class. This proves the relation between (1) and (3).
Thus the three classes are related functorially: realization carries the mixed-Hodge-module extension to the perverse extension, while hypercohomology carries it to the LMHS extension class. Finally, all three descriptions arise from the same nearby-cycle and variation data because $\mathcal{P}^{H}$ is identified in Proposition 6.10 with the cone of the global variation morphism, so both its realization and its induced hypercohomological extension inherit that same origin. ∎
@@ 8.4. Toward the domain-wall interpretation
The significance of the present paper is that this interpretation no longer rests only on the perverse side. Theorems 6.9 and 7.7 show that the same localized quotient also contributes the vanishing part of the limiting mixed Hodge structure on hypercohomology. Thus the bulk/localized-sector interpretation is simultaneously visible in the mixed-Hodge-module extension, in its perverse realization, and in the Hodge-theoretic degeneration data. This is the precise mathematical content that later physical applications may safely use.
@@ 9.1. What has been proved
The main result of the paper is the construction of a mixed Hodge module
@@
whose realization is the corrected perverse object associated with a finite multi-node conifold degeneration. Its singular quotient is the finite point-supported object
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1),$
and it fits into the corresponding exact sequence with $IC^{H}_{X_{0}}$.
The paper also proves that this quotient is the same object that appears as the finite local vanishing sector in the nearby-cycle formalism on the Hodge-theoretic side. In particular, the corrected extension is compared at three levels: as an extension in mixed Hodge modules, as its realized extension in perverse sheaves, and as an induced extension on hypercohomology in the category of mixed Hodge structures.
Finally, the finite multi-node setting is shown to admit a nodewise organization of the relevant extension spaces, both in mixed Hodge modules and on the perverse side. This gives the extension data a form that is suitable for the later structural questions discussed below.
@@ 9.2. Consequences for subsequent work
The results proved here provide the mixed-Hodge-module framework needed for several later directions.
First, the object
@@
makes it possible to formulate subsequent questions directly in mixed Hodge modules rather than only through realization or through nearby-cycle heuristics.
Second, the quotient
$\bigoplus_{k=1}^{r}i_{k*}\mathbb{Q}^{H}_{\{p_{k}\}}(-1)$
now has three compatible roles: it is the singular quotient in $MHM(X_{0})$, its realization is the finite node-supported quotient on the perverse side, and on hypercohomology it identifies the finite local vanishing sector in the limiting mixed Hodge structure.
Third, the nodewise decomposition of the extension space gives the finite multi-node setting a precise extension-theoretic organization. This is the structural input for later work on stronger comparison theorems, nodewise generators, and possible quiver- or schober-type refinements.
In this sense, the present paper supplies the mixed-Hodge-module object and comparison framework on which those later developments can build.